OpenStudy (anonymous):

Two groups of tourists each have 60 people. If 3/4 of the first group and 2/3 of the second group board buses to travel to a museum, how many more people in the first group board buses than in the second group?

OpenStudy (anonymous):

Try to work out how many people are in each bus group. How many people is 3/4 of 60?

OpenStudy (anonymous):

3/4=9/12 2/3=8/12 1/12 more than the other

OpenStudy (anonymous):

60/12=5 more people

OpenStudy (gw2011):

First Group: (3/4)(60)=45 Second Group: (2/3)(60)=40 45-40=5 more people board buses in the first group